OpenStudy (anonymous):

Your friend is trying to calculate the height of a nearby oak tree. You tell him that you learned how to use similar triangles in Geometry class. You tell your friend to measure his height (75 inches) and you measure the length of his shadow (48 inches). Both of you measure the length of the tree’s shadow (38 feet). How tall is the tree (in feet)? Round to the nearest hundredth. help please!!1

OpenStudy (bahrom7893):

75/48 = x/38

OpenStudy (bahrom7893):

x is the height of the tree. The ratios stay the same (similar triangles)

OpenStudy (bahrom7893):

woah wait.. 38 feet, convert that to inches

OpenStudy (bahrom7893):

38ft = 456 inches

OpenStudy (bahrom7893):

75/48 = x/456 x = 456*75/48

OpenStudy (anonymous):

Well you have to prove that they are similar triangles before using any of this.

OpenStudy (bahrom7893):

x = 712.5 inches. FFM, read the question...

OpenStudy (bahrom7893):

oh in feet.. hold on convert that back to feet

OpenStudy (anonymous):

I have.

OpenStudy (bahrom7893):

59.38 feet is the answer

OpenStudy (anonymous):

There is now way we could write "75/48 = x/38" without establishing the corresponding similar triangles.

OpenStudy (bahrom7893):

Listen all shadow problems always, ALWAYS have to do with similar triangles. There's no way you cannot be perpendicular to your shade on the ground, that means your height and the length of your shadow is always going to form a right triangle (same goes for any other object). Also, we're assuming that you don't measure your shadow at noon and the shadow of the tree early in the morning. And we're assuming that both the tree and you are perpendicular to the ground, and not at some random angle, so yea I think we can write that.

OpenStudy (anonymous):

>> Also, we're assuming that you don't measure your shadow at noon and the shadow of the tree early in the morning. It's actually assumed that the calculation is done simultaneously, at the same time, (to get a good answer, recall the position of earth and sun changes (slowly) during the whole day. >>so yea I think we can write that. yes you can, and it will give you the correct answer in every case but it's not at all insightful and there is a good chance of making mistakes of the kinds like you did here. The more you work the more you learn :)

OpenStudy (bahrom7893):

ffm go troll somewhere else and stop trying to teach me geometry. And I did not make any mistakes on here, so if you see any, you can go back and review.

OpenStudy (anonymous):

Apart from the fact your answer is not insightful there is also a mistake: \( 75/48 = x/38 \) they are not in same units so it's incorrect And please be respectful!